The function x(r) = _ 2.17e-1.19, r(t)--2.17e-1191 describes the position, in meters, of a particle on the x-axis as a function of time t, in seconds. Find the particles velocity at time t- 3.17 s Number m/ s
Answer #1

Solution:

Given x(t) = -2.17 * e^(-1.19t)

=> v(t) = dx(t)/dt

=> v(t) = -2.17 * e^(-1.19t) * -1.19

=> v(t) = 2.5823 * e^(-1.19t)

=> v(3.17) = 2.5823 * e^(-1.19 * 3.17)

=> v(3.17 ) = 0.0594 m/s

thanks.